There are 15 girls and 10 boys in a class. The teacher randomly selects one student to go to the library and then another student to go to the library. What is the probability that both students randomly picked are both girls?
Answer by checkley75(3666)

this solution assumes there are no replacements:
picking the first girl is 15/25=3/5
the second pick is 14/24=7/12
now we multiply these two fractions
3/5*7/12=21/60=7/20=.35 is the probability that the first 2 picks are both girls.